OpenStudy (anonymous):

Find the equation of the quadratic function with zeros at -1, and 1 and vertex at (0, -6). Find the equation of the quadratic function with zeros 10 and 14 and vertex at (12, -8).

OpenStudy (amistre64):

first you construct it for the zeros: we know the xs so find the "a" "b" to fit (-1+a)(1+b) = 0 a = 1 and b = -1 (x+1)(x-1) = 0 now, when x=0, what do we have to multiply it by to get it to equal -6?

OpenStudy (amistre64):

n(0+1)(0-1) = -6 n(1)(-1) = -6 -n = -6 n=6 6(x-1)(x+1)=0 should suffice

OpenStudy (amistre64):

there might be an easier way to do that, but thats just what came first to me head :)

OpenStudy (anonymous):

and what about the second one?

OpenStudy (amistre64):

y = a(x-h)^2 + k where vertex = (h,k)

OpenStudy (amistre64):

the second one is the same process; just different numbers

OpenStudy (anonymous):

y = 2x2 + 48x + 280 y = -2x2 - 48x + 280 y = -2x2 + 48x - 280 y = 2x2 - 48x + 280 these are my choices

OpenStudy (amistre64):

since we know 12,-8 y = a(x-h)^2 + k y = a(x-12)^2 -8 is one setup

OpenStudy (amistre64):

i like the other method better tho

OpenStudy (anonymous):

http://i729.photobucket.com/albums/ww292/akatushi101/A2_Image_SemesterAFinal_278.gif Given the following triangle, solve for angle C. is it 56 degrees?